OpenStudy (anonymous):

Math Analysis: Convert the radian measure to degree measure. Use the value of pie found on your calculator and round answers to two decimal places.

OpenStudy (anonymous):

1.345

OpenStudy (anonymous):

pie=180

OpenStudy (anonymous):

okay, so how do i do this

OpenStudy (anonymous):

77.0628234 degree = 1.345 rad

OpenStudy (anonymous):

how did u get that on the calculator

OpenStudy (anonymous):

Convert by multiplying by 180/pie

OpenStudy (anonymous):

ok so i get 57.2 on my calculator

OpenStudy (anonymous):

In Degree mode, the r symbol can be used to convert radians to degrees. The r is found 2nd APPS(Angle) #3 r.

OpenStudy (anonymous):

so i got 57.2 on my calculator how do i get the answer lol

OpenStudy (anonymous):

1 radian is 180/pi degrees and 1 degree is pi/180 radians just multiply that to get wot u need

OpenStudy (anonymous):

or type in google search box

OpenStudy (anonymous):

so i divide 1.345/180 in radians on my calculator?

OpenStudy (anonymous):

1.345*(22/7/180)

OpenStudy (anonymous):

on a ti-83 or ti-84, all you need to do is press the "mode" button. this takes you to a screen where you can choose whether to be in radians or degrees.

OpenStudy (anonymous):

im on degrees right now

OpenStudy (anonymous):

go to radians

OpenStudy (anonymous):

ok, i got .02348

OpenStudy (anonymous):

on a ti-89 titanium calculator goto Mode, then go to Angle and right arrow click to whichever unit you want.

OpenStudy (anonymous):

i only have a TI-30XIIS

OpenStudy (anonymous):

just go to calculation mode (simple multiplication addition mode) then enter rad you want to convert and multiply it by (180/22/7

OpenStudy (anonymous):

i am, im puting this in my calculator: 1.345 x (180/22/7) in degrees and i'm getting 1.57

OpenStudy (anonymous):

To convert radians to degrees, we make use of the fact that p radians equals one half circle, or 180º. This means that if we divide radians by p, the answer is the number of half circles. Multiplying this by 180º will tell us the answer in degrees. So, to convert radians to degrees, multiply by 180/p,

OpenStudy (anonymous):

so, 1.345 x (180/ what)

OpenStudy (anonymous):

p=pie

OpenStudy (anonymous):

1.345 x (180/ pie) in degrees = 0

OpenStudy (anonymous):

don't go in degrees mode

OpenStudy (anonymous):

http://mathbits.com/mathbits/tisection/trig/degrees.htm go to this site

OpenStudy (anonymous):

1.345 x (180/ pie) in radians = 0

OpenStudy (anonymous):

got o the site please it shows converting rad to deg & deg to rad with calculator images

OpenStudy (anonymous):

go to

OpenStudy (anonymous):

im following what you were telling me and it's not working

OpenStudy (diyadiya):

Pi = 3.14159265................

OpenStudy (anonymous):

i know

OpenStudy (anonymous):

Convert the radian measure to degree measure. Use the value of pie found on your calculator and round answers to two decimal places. 1.345

OpenStudy (anonymous):

Method 1: (in Radian or Degree mode) Convert by multiplying by . 180/pi Mode set to Degree Method 2: (in Degree mode) In Degree mode, the r symbol can be used to convert radians to degrees. The r is found 2nd APPS(Angle) #3 r. With the mode set to Degree, type (3pi2)^r

Parth (parthkohli):

\( \color{Black}{\Rightarrow 2\pi = 360^{\circ} }\) You can use this and find any measure in radians.

OpenStudy (anonymous):

I told it many time but she/he is not understanding

Parth (parthkohli):

I believe that \(1.34 = \pi \div 3\) Use the equation ajay and I gave to find \(\pi \div 3\)

OpenStudy (anonymous):

ok i got 57.30 when i divided 180/pie

OpenStudy (anonymous):

it says use calculator

OpenStudy (anonymous):

see not understanding

Parth (parthkohli):

\( \color{Black}{\Rightarrow 2\pi :360 ={\Large{\pi \over 3}}:x }\)

OpenStudy (anonymous):

2pie/360=180

OpenStudy (anonymous):

it should be 1

OpenStudy (anonymous):

no 180

OpenStudy (anonymous):

i read the online article and i still don't get it

OpenStudy (anonymous):

what is 2nd APPS(Angle) #3 r.

OpenStudy (anonymous):

Radians to degrees The general formula for converting from degrees to radians is to simply multiply the number of degree by 180°/(Π) Convert 1.4 radians into degrees: 1.4 (180°/Π) = 80.2 °

OpenStudy (anonymous):

nvm i found it out myself :P

OpenStudy (anonymous):

video http://www.youtube.com/watch?v=KiP10rUzDP4